OpenStudy (anonymous):

how many solutions does the equation sin(4x) = 1/2 have on the interval (0,2pi)

OpenStudy (anonymous):

count intersection blue and red

OpenStudy (anonymous):

this only one way to solve the question..

OpenStudy (anonymous):

say 4x=a sina=1/2 what could be a? a=pi/6 right what else? a=n*pi+pi/6 n in Z 4x=n*pi+pi/6 x=1/4(n*pi+pi/6) n=0,1,2,3,4,5,6,7,8 but not 9 because n=9 1/4(9*pi+pi/6=7.19>2pi=6.28

OpenStudy (anonymous):

sorry n=8 is also not working 1/4(8*pi+pi/6=6.41>2pi=6.28 so answer is 8

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

np